# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ Abstract
Abstract. We give the first exact Bayesian methodology for the problem of inference in discretely observed regime switching diffusions. We design an MCMC and an MCEM algorithm that target the exact posterior of diffusion parameters and the latent regime process. The algorithms are exact in the sense that they target the correct posterior distribution of the continuous model, so that the errors are due to Monte Carlo only. Switching diffusion models extend ordinary diffusions by allowing for jumps in instantaneous drift and volatility. The jumps are driven by a latent, continuous time Markov switching process. We illustrate the method on numerical examples, including an empirical analysis of the method’s scalability in the length of the time series, and find that it is comparable in computational cost with discrete approximations while avoiding their shortcomings.
@@ 1 Introduction
Many stochastic phenomena are modelled by an observable process whose parameters depend on a time-changing unobserved regime, commonly modelled as a finite state-space Markov process. If the model allows for serial dependence after conditioning on the regime, we speak of a Markov switching model. Such models are most common in economics and finance, where they were first proposed to infer business cycles from GDP growth data (hamilton1989new). Other economic time series exhibiting cyclical regime shifts include exchange rates (engel1990long), interest rates (cai1994markov), stock prices (hamilton1994autoregressive), commodity prices (fong2002markov) and energy prices (mount2006predicting). Regime switching processes also lend themselves to modelling structural breaks in economic regimes, such as in kim1999has; mcconnell2000output. While discrete time models are dominant in econometrics, Markov switching models also have a natural continuous time formulation as Markov switching diffusions, i.e. diffusion processes whose drift and volatility functions change according to a continuous time Markov jump process. For example, yan2014moving apply a driftless switching model to the task of animal tracking. Moreover, mathematicians have long investigated stability and optimal control of such models (ghosh1993optimal; basak1996stability; mao2006stochastic). Conversely, since the transition law of most diffusion models is not analytically available, likelihood-based inference is not immediately possible.
@@
As exemplified by hamilton1989new, most of the existing literature on Markov switching uses discrete time models, thereby avoiding the technical challenges of the continuous time setting. Others seek to address the problem by either restricting the diffusion process to an analytically tractable family (blackwell2003bayesian; hahn2010markov), or by using an Euler-Maruyama-type discrete time approximation of the process dynamics (hodgson1999bayesian; liechty2001markov). Higher-order approximation schemes were proposed by ait2002maximum and applied to Markov switching diffusions in choi2009regime. Since such approximations are biased, an MCMC algorithm that relies on them yields samples from an approximate posterior. As the Euler-type discretization is refined, the approximation to the posterior improves consistently. Nonetheless, whereas Monte Carlo error control is well-studied and understood, the discretization bias resulting from Euler approximation is difficult to quantify in any given finite sample setting, and more so in the presence of regime switching. Asymptotically and in the instance of Ito diffusions, the error of Euler-type schemes in estimating test functions is of order $\mathcal{O}(\text{effort}^{-1/3})$ (duffie1995efficient), which in special cases can be reduced to $\mathcal{O}(\text{effort}^{-1/2})$ (giles2008multilevel). Accordingly, wherever feasible, we deem it preferable in terms of asymptotic efficiency and transparency of error to apply exact methods, understood as being subject to Monte Carlo error only. This applies in particular to the Markov switching context, where approximation error is less understood, see e.g. mao2007approximations, and which presents particular problems due to off-equilibrium effects upon a change of latent regime, where drift can be large and highly variable. In addition, by preserving the full continuous-time setting, our solution provides exact inference on both the latent and the observable process along their entire continuous domain - see Figure 1, which pertains to the animal tracking application explored in Section 5. In particular, we observe that predictive intervals automatically account for periods of low (“resting”) and high variability (“moving”) of observations. In the same section, we also empirically examine the extent of the bias when estimating the tracking model with an approximate algorithm.
@@
In that light, we will construct a Markov Chain with stationary distribution corresponding to the exact posterior. As a result, posterior summaries are subject to Monte Carlo estimation error only, and if the Monte Carlo Central Limit Theorem holds, we recover $\mathcal{O}(\text{effort}^{-1/2})$ asymptotics. This leverages an extensive literature on exact simulation of Ito diffusions (e.g. beskos2008factorisation), as well as Bayesian posterior simulation for Ito diffusions (gonccalves2017barker) and jump diffusions (gonccalves2023exact). Those methods incorporate the missing, infinite-dimensional diffusion paths in order to exploit the complete diffusion likelihood. Critically, even though the algorithm targets a distribution on the space of infinite-dimensional paths, it only requires the evaluation of the diffusion path on a finite subset of times, which is extended as required for the propagation of the algorithm by interpolating the previously revealed path skeleton. This is known as retrospective simulation. While various retrospective techniques have been applied to construct exact MCMC algorithms, we focus on the Bernoulli factory MCMC approach, seen e.g. in gonccalves2017barker; gonccalves2023exact, which allows for the implementation of accept/reject coin flips without explicitly evaluating the (potentially inctractable) probability of acceptance. Relative to other approaches to constructing exact MCMC algorithms, such as pseudo-marginal MCMC, this has the benefit of minimizing the extent of data augmentation, which is liable to negatively impact Markov chain mixing and inflate estimation variance, sometimes in intransparent ways, see e.g. andrieu2009pseudo. Indeed, Bernoulli MCMC algorithms also have more transparent failure modes, in that their iteration time will inflate, unlike the deterioration in mixing that occurs e.g. in pseudo-marginal algorithms (andrieu2009pseudo). Moreover, in our experiments, we observe close to linear scaling of computational cost in data size, as opposed to pseudo-marginal algorithms, which are usually quadratic in data size (schmon2021large).
Though we adopt much of the framework seen in gonccalves2023exact and leverage some innovations from vats2020efficient; stumpf2024bf, we surmount a range of challenges particular to the Markov switching setting, and in doing so improve on established exact methods. Those challenges arise due to the presence of the latent Markov jump process, which can result in more erratic drift and volatility patterns. It is only through careful exploitation of the model structure, flexible initialization and adaptation procedures, as well as improvements to the algorithmic complexity of Bernoulli MCMC methods, that practical MCMC algorithms can arise in this context, even for modest sample sizes. Indeed, our results demonstrate good scalability properties without any ad-hoc tuning. In doing so, we set a template for how to apply exact methods in complex latent process models.
@@
The paper is organized as follows. Section 1 continues with a presentation of the model of interest and the notational conventions. Section 2 introduces the augmentation scheme that underlies our inference algorithms. Section 3 describes an MCMC algorithm for posterior sampling, and elaborates on various underlying techniques. Section 4 describes an analogous MCEM algorithm for MAP estimation. Section 5 demonstrates both posterior sampling and point estimation as well as approximation bias on an authentic animal tracking time series. We close with a study of the algorithm’s scalability in Section 6.
@@ 1.1 The Markov Switching Diffusion Setting
The regime-switching framework in this paper is as follows. Let $Y$ be a latent, discrete space, continuous time Markov jump process with regimes $\mathcal{Y}=\{1,\dots,k\}$, taking values in the set $\mathcal{K}$ of $([0,\omega]\mapsto\mathcal{Y})$ Càdlàg functions. The jump process evolves according to its generator matrix $\lambda$, where $\lambda_{i,j\neq i}\geq 0$ are the jump rates from regime $i$ to $j$ and the diagonal elements are given by $\lambda_{ii}=-\sum_{i\neq j}\lambda_{ij}$. We follow the common convention of denoting the exit rates $\lambda_{i}=-\lambda_{ii}$. The density function of $Y$ is defined with respect to the measure $\mathbb{L}$ induced by a rate 1 marked Poisson process. Define the Markov switching stochastic differential equation (SDE)
@@
where $W$ is a standard Brownian motion and $\theta$ is a parameter vector. The methodology in this paper naturally extends to multivariate diffusions, with the caveat of stricter requirements on the functional form of $\sigma_{\theta}(V_{t})$. Suppose that the SDE admits a unique solution for every $y\in\mathcal{K}$ and $\theta$ and therefore a Markov transition density $\pi(v_{t+\epsilon}|v_{t},y,\theta)$. Assume that $V$ with state space $\mathcal{V}$ is observed at times $0,s_{1},s_{2},\dots$ with values $v_{0},v_{s_{1}},v_{s_{2}},\dots$ while all other quantities are unknown, i.e. $\theta$ and $\lambda$ denote realizations of the random variables $\Theta$ and $\Lambda$. Our aim is to generate samples from the exact posterior $\pi(y,\theta,\lambda|v_{0},v_{s_{1}},v_{s_{2}},\dots)$ for a given product prior $\pi(\theta,\lambda)=\pi(\theta)\pi(\lambda)$. The factorization of the volatility term need not be unique and this arbitrariness does not affect the algorithm presented in the paper. Notice that for path functionals $g(v)$, the method also allows for estimation of posterior expectations $\operatorname{\mathrm{E}}\left[g(V)|v_{0},v_{s_{1}},v_{s_{2}},\dots\right]$. The overarching goal is to devise a Markov chain Monte Carlo algorithm that targets the posterior
$\pi(y,\theta,\lambda|v_{0},v_{s_{1}},v_{s_{2}},\dots)\propto\pi(y|\lambda)\pi(\theta)\pi(\lambda)\pi(v_{s_{1}}|v_{0},y,\theta)\pi(v_{s_{2}}|v_{s_{1}},y,\theta)\cdots.$ (2)
Except for rare special cases, the conditional transition density $\pi(v_{t+\epsilon}|v_{t},y,\theta)$ is intractable. Hence, standard methods of likelihood-based inference are not directly applicable. Therein lies the fundamental challenge to inference in continuous time models.
@@ 1.2 Conventions
Throughout the paper, random variables will be written in uppercase letters and realizations thereof in lower case. Moreover, for a random variable $A$ and realization $a$, $\pi(a)$ refers to the density at $a$ if $A$ is continuous and the probability mass at $a$ if $A$ is discrete. For a probability measure $\mathbb{M}$ on $(A,B)$, $\mathbb{M}_{|a}$ denotes the conditional measure on $B$ after observing $\{A=a\}$. For three sets $a,b,c$ such that $a\subseteq b$ and a function $f:b\to c$, $f(a)$ denotes $\{f(\dot{a}):\dot{a}\in a\}$. Analously, given a continuous path $v$ and a set of times $s$, we use the notation $v_{s}$ to refer to $\{v_{s}:\dot{s}\in s\}$. $\{(\dot{s}\sim\ddot{s})\in s\}$ refers to the set of neighboring pairs in $s$, and we use $\omega$ to denote the “end of time”.
@@ 2 Data Augmentation Strategy
In this section, we derive the complete transition density of the model with respect to an appropriate dominating measure, amenable to efficient Gibbs sampling. Since we presume that the transition density is intractable, it is not possible to target the marginal posterior over $(Y,\Theta,\Lambda)$ with standard methods. We follow the common strategy of augmenting the state space with the missing diffusion bridges $V_{(\dot{s},\ddot{s})}$ in between the observation pairs $(\dot{s}\sim\ddot{s})\in s$. This results in a posterior distribution with convenient conditional independence structure. Indeed, it is useful to segment the diffusion path according to the union of observations $V_{s}$ and imputed values $V_{R}=v_{r}$, where $R$ is the set of times where $Y$ changes values. Defining $\tau=s\cup r$, this yields the complete likelihood
@@
with respect to an appropriate dominating measure. Such a centered parameterization is not amenable to Gibbs sampling because the full conditionals $\pi(v_{[0,\omega]\mathbin{\backslash}s}|v_{s},y,\theta)$ and $\pi(\theta|v,y)$ are mutually singular for distinct values $\theta$ and $\theta^{\dagger}$, as expanded upon in roberts2001inference. Indeed, the quadratic variation increment $\mathop{}\!\mathrm{d}\langle V\rangle_{t}=\sigma_{\theta}^{2}(V_{t})\rho_{\theta}^{2}(Y_{t})\mathop{}\!\mathrm{d}t$ is a deterministic function of $V$, $Y$ and $\theta$, so $\pi(\theta|v,y)$ can only assign positive probability to values of $\theta$ for which $\langle V\rangle_{t}$ is preserved. The standard solution consists of changing variables to a non-centered parameterization, where the diffusion bridges and $\theta$ are a priori independent. Equivalently, the change of variables results in a density with respect to a fixed dominating measure. To carry out this change of variables, we define the Lamperti transform
@@
$\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\rho_{\theta}(y_{\dot{\tau}})z_{t}+\eta_{\theta}(v_{\dot{\tau}})+(\eta_{\theta}(v_{\ddot{\tau}})-\eta_{\theta}(v_{\dot{\tau}}))\frac{t-\dot{\tau}}{\ddot{\tau}-\dot{\tau}}.\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (6)
@@
  • The Novikov condition applies, i.e. $\operatorname{\mathrm{E}}_{X_{(0,\omega]}}\left[\exp\left\{\int_{0}^{\omega}\delta_{\theta}^{2}(X_{t},b)\mathop{}\!\mathrm{d}t\right\}|\{Y_{(0,\omega]}=b\},x_{0},\theta\right]<\infty$ for every $\theta,b\in\mathcal{Y},x_{0}\in\eta_{\theta}(\mathcal{V}),\omega<\infty$. This is sufficient, albeit not necessary.
@@
$\varphi_{\theta}(a,b)=\frac{1}{2}\left\lparen\frac{\delta_{\theta}^{2}(a,b)}{\rho_{\theta}^{2}(b)}+\partial_{a}\delta_{\theta}(a,b)\right\rparen,$ (7)
@@
$h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\lvert\eta_{\theta}^{\prime}(v_{\ddot{\tau}})\rvert\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]e^{\rho_{\theta}^{-2}(y_{\dot{\tau}})\{\Delta_{\theta}(\eta_{\theta}(v_{\ddot{\tau}}),y_{\dot{\tau}})-\Delta_{\theta}(\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}})\}}.$ (9)
@@
$\begin{split}\pi(v_{\tau\mathbin{\backslash}\{0\}},z|v_{0},y,\theta)=\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau}\pi(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)\end{split}$ (12)
marginalizes to $\pi(v_{s\mathbin{\backslash}\{0\}}|v_{0},y,\theta)$ by Proposition 1, and its dominating measure is suitably invariant in $\theta$. Note that $\pi(v_{\tau\mathbin{\backslash}\{0\}},z|v_{0},y,\theta)$ cannot be directly evaluated due to the presence of a path integral over $z$. Indeed, $z$ cannot even be exhaustively stored in memory. Nonetheless, the algorithm presented in the next section is for all intents and purposes an MCMC algorithm on an infinite-dimensional state space, so we mostly treat the actual representation of $z$ as an implementation detail, and our notation refers to the full path $z$, unless that implementation is relevant.

### v2

@@ Abstract
We develop the first exact Bayesian methodology for the problem of inference in discretely observed regime switching diffusions. Switching diffusion models extend ordinary diffusions by allowing for jumps in instantaneous drift and volatility. The jumps are driven by a latent, continuous time Markov switching process. We address the problem through an MCMC and an MCEM algorithm that target the exact posterior of diffusion parameters and the latent regime process. The algorithms are exact in the sense that they target the correct posterior distribution of the continuous model, so that the errors are due to Monte Carlo only. We illustrate the method on numerical examples, including an empirical analysis of the method’s scalability in the length of the time series, and find that it is comparable in computational cost with discrete approximations while avoiding their shortcomings.
Code: https://github.com/timsf/markov-switch
Keywords: diffusions, time series, regime switching, intractable likelihood, Bernoulli MCMC, Markov chain Monte Carlo.
MSC2020 subject classifications: Primary 62-08; Secondary 62F15, 62M05
@@ 1 Introduction
Many stochastic phenomena are modelled by an observable process whose parameters depend on a time-changing unobserved regime, commonly modelled as a finite state-space Markov process. If the model allows for serial dependence after conditioning on the regime, we speak of a Markov switching model. Such models are most common in economics and finance, where they were first proposed to infer business cycles from GDP growth data [24]. Many other economic time series exhibit cyclical regime shifts, such as exchange rates [16], interest rates [10], stock prices [25], commodity prices [17] and energy prices [36]. Regime switching processes also lend themselves to modelling structural breaks in economic regimes, such as in [28, 34]. While discrete time models are dominant in econometrics, Markov switching models also have a natural continuous time formulation as Markov switching diffusions, i.e. diffusion processes whose drift and volatility functions change according to a continuous time Markov jump process. For example, [46] apply a driftless switching model to the task of animal tracking. Moreover, mathematicians have long investigated stability and optimal control of such models [19, 4, 32]. Conversely, since the transition law of most diffusion models is not analytically available, likelihood-based inference is not immediately possible.
@@
As exemplified by [24], most of the existing literature on Markov switching uses discrete time models, thereby avoiding the technical challenges of the continuous time setting. Others seek to address the problem by either restricting the diffusion process to an analytically tractable family [9, 23], or by using an Euler-Maruyama-type discrete time approximation of the process dynamics [27, 31]. Higher-order approximation schemes were proposed by [1] and applied to Markov switching diffusions in [12]. Since such approximations are biased, an MCMC algorithm that relies on them yields samples from an approximate posterior. As the Euler-type discretization is refined, the approximation to the posterior improves consistently. Nonetheless, whereas Monte Carlo error control is well-studied and understood, the discretization bias resulting from Euler approximation is difficult to quantify in any given finite sample setting, and more so in the presence of regime switching. Asymptotically and in the instance of Ito diffusions, the error of Euler-type schemes in estimating test functions is of order $\mathcal{O}(\text{effort}^{-1/3})$ [15], which in special cases can be reduced to $\mathcal{O}(\text{effort}^{-1/2})$ [20]. Accordingly, wherever feasible, we deem it preferable in terms of asymptotic efficiency and transparency to apply exact methods, understood as being subject to Monte Carlo error only. This applies in particular to the Markov switching context, where approximation error is less understood, see e.g. [33], and which presents particular problems due to off-equilibrium effects upon a change of latent regime, where drift can be large and highly variable. In addition, by preserving the full continuous-time setting, our solution provides exact inference for both the latent and the observable process along their entire continuous domain - see Figure 1, which pertains to the animal tracking application explored in Section 5. In particular, we observe that predictive intervals automatically account for periods of low (“resting”) and high variability (“moving”) of observations. In the same section, we also empirically examine the extent of the bias when estimating the tracking model with an approximate algorithm.
@@
In that light, we will construct a Markov Chain with stationary distribution corresponding to the exact posterior. As a result, posterior summaries are subject to Monte Carlo estimation error only, and if the Monte Carlo Central Limit Theorem holds, we recover $\mathcal{O}(\text{effort}^{-1/2})$ asymptotics. This leverages an extensive literature on exact simulation of Ito diffusions [5] as well as Bayesian posterior simulation for Ito diffusions [21] and jump diffusions [22]. Those methods incorporate the missing, infinite-dimensional diffusion paths in order to exploit the complete diffusion likelihood. Critically, even though the algorithm targets a distribution on the space of infinite-dimensional paths, it only requires the evaluation of the diffusion path on a finite subset of times, which is extended as required for the propagation of the algorithm by interpolating the previously revealed path skeleton. This is known as retrospective simulation. While various retrospective techniques have been applied to construct exact MCMC algorithms, we focus on the Bernoulli factory MCMC approach, seen e.g. in [21, 22], which allows for the implementation of accept/reject coin flips without explicitly evaluating the potentially intractable probability of acceptance. Relative to other approaches to constructing exact MCMC algorithms, such as pseudo-marginal MCMC, this has the benefit of minimizing the extent of data augmentation, which is liable to negatively impact Markov chain mixing and inflate estimation variance, sometimes in intransparent ways. As it happens, Bernoulli MCMC algorithms also have more transparent failure modes, in that their iteration time will inflate, unlike the deterioration in mixing that occurs e.g. in pseudo-marginal algorithms [42, 2]. Moreover, in our experiments, we observe close to linear scaling of computational cost in data size. This stands in contract with pseudo-marginal algorithms, where a quadratic cost per iteration has to be incurred in order to keep the variance of the likelihood estimator constant as data size increases, as recommended e.g. by [42, 14].
Though we adopt much of the framework seen in [22] and leverage some innovations from [45, 44], we surmount a range of challenges particular to the Markov switching setting, and in doing so improve on established exact methods. Those challenges arise due to the presence of the latent Markov jump process, which can result in more erratic drift and volatility patterns. It is only through careful exploitation of the model structure, flexible initialization and adaptation procedures, as well as improvements to the algorithmic complexity of Bernoulli MCMC methods, that practical MCMC algorithms can arise in this context, even for modest sample sizes. Indeed, our results demonstrate good scalability properties without any ad-hoc tuning. In doing so, we set a template for how to apply exact methods in complex latent process models.
@@
The paper is organized as follows. Section 1 continues with a presentation of the model of interest and the notational conventions. Section 2 introduces the augmentation scheme that underpins our inference algorithms. Section 3 describes an MCMC algorithm for posterior sampling, and elaborates on various underlying techniques. Section 4 describes an analogous MCEM algorithm for MAP estimation. Section 5 demonstrates both posterior sampling and point estimation as well as approximation bias on an authentic animal tracking time series. We close with a study of the algorithm’s scalability in Section 6.
@@ 1.1 The Markov Switching Diffusion Setting
The regime-switching framework in this paper is as follows. Let $Y$ be a latent, discrete space, continuous time Markov jump process with regimes $\mathcal{Y}=\{1,\dots,k\}$, taking values in the set $\mathcal{K}$ of $([0,\omega]\mapsto\mathcal{Y})$ Càdlàg functions. The jump process evolves according to its generator matrix $\lambda$, where $\lambda_{i,j\neq i}\geq 0$ are the jump rates from regime $i$ to $j$ and the diagonal elements are given by $\lambda_{ii}=-\sum_{i\neq j}\lambda_{ij}$. We follow the convention of denoting the exit rates $\lambda_{i}=-\lambda_{ii}$. The density function of $Y$ is defined with respect to the measure $\mathbb{L}$ induced by a rate 1 marked Poisson process. Define the Markov switching stochastic differential equation (SDE)
@@
where $W$ is a standard Brownian motion and $\theta$ is a parameter vector. The methodology in this paper naturally extends to multivariate diffusions, with the caveat of stricter requirements on the functional form of $\sigma_{\theta}(V_{t})$. Suppose that the SDE admits a unique solution for every $y\in\mathcal{K}$ and $\theta$ and therefore a Markov transition density $\pi(v_{t+\epsilon}|v_{t},y,\theta)$. Assume that $V$ with state space $\mathcal{V}$ is observed at times $0,s_{1},s_{2},\dots$ with values $v_{0},v_{s_{1}},v_{s_{2}},\dots$ while all other quantities are unknown, i.e. $\theta$ and $\lambda$ denote realizations of the random variables $\Theta$ and $\Lambda$. The factorization of the volatility term need not be unique and this arbitrariness does not affect the algorithm presented in the paper. The overarching goal is to devise a Markov chain Monte Carlo algorithm that targets the posterior
$\pi(y,\theta,\lambda|v_{0},v_{s_{1}},v_{s_{2}},\dots)\propto\pi(y|\lambda)\pi(\theta)\pi(\lambda)\pi(v_{s_{1}}|v_{0},y,\theta)\pi(v_{s_{2}}|v_{s_{1}},y,\theta)\cdots$ (2)
for a given product prior $\pi(\theta,\lambda)=\pi(\theta)\pi(\lambda)$. Except for rare special cases, the conditional transition density $\pi(v_{t+\epsilon}|v_{t},y,\theta)$ is intractable. Hence, standard methods of likelihood-based inference are not directly applicable. Therein lies the fundamental challenge to inference in continuous time models.
@@ 1.2 Conventions
Throughout the paper, random variables will be written in uppercase letters and realizations thereof in lower case. Moreover, for a random variable $A$ and realization $a$, $\pi(a)$ refers to the density at $a$ if $A$ is continuous and the probability mass at $a$ if $A$ is discrete. For a probability measure $\mathbb{M}$ on $(A,B)$, $\mathbb{M}_{|a}$ denotes the conditional measure on $B$ after observing $\{A=a\}$. For three sets $a,b,c$ such that $a\subseteq b$ and a function $f:b\to c$, $f(a)$ denotes $\{f(\dot{a}):\dot{a}\in a\}$. Analogously, given a continuous path $v$ and a set of times $s$, we use the notation $v_{s}$ to refer to $\{v_{s}:\dot{s}\in s\}$. For an ordered set $s=\{s_{0}<s_{1}<\dots\}$, $\{(\dot{s}\sim\ddot{s})\in s\}$ refers to the set of paired neighbors $(s_{0},s_{1}),(s_{1},s_{2}),\dots$, and we use $\omega$ to denote the “end of time”.
@@ 2 Data Augmentation Strategy
In this section, we derive the complete transition density of the model with respect to an appropriate dominating measure, amenable to efficient Gibbs sampling. Since we presume that the transition density is intractable, it is not possible to target the marginal posterior over $(Y,\Theta,\Lambda)$ with standard methods. We follow the common strategy of augmenting the state space with the missing diffusion bridges $V_{(\dot{s},\ddot{s})}$ in between the observation pairs $(\dot{s}\sim\ddot{s})\in s$. This results in a posterior distribution with convenient conditional independence structure. In fact, it is useful to segment the diffusion path according to the union of observations $V_{s}$ and imputed values $V_{R}=v_{r}$, where $R$ is the set of times where $Y$ changes values. Defining $\tau=s\cup r$, this yields the complete likelihood
@@
[…]
@@
$\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})=\eta_{\theta}(v_{\dot{\tau}})+(\eta_{\theta}(v_{\ddot{\tau}})-\eta_{\theta}(v_{\dot{\tau}}))\frac{t-\dot{\tau}}{\ddot{\tau}-\dot{\tau}}+\rho_{\theta}(y_{\dot{\tau}})z_{t}.\qquad(t\in(\dot{\tau},\ddot{\tau}))$ (6)
@@
  • The Novikov condition applies, i.e. $\operatorname{\mathrm{E}}_{X_{(0,\omega]}}\left[\exp\left\{\int_{0}^{\omega}\delta_{\theta}^{2}(X_{t},b)\mathop{}\!\mathrm{d}t\right\}|\{Y_{(0,\omega]}=b\},x_{0},\theta\right]<\infty$ for every $\theta,b\in\mathcal{Y},x_{0}\in\eta_{\theta}(\mathcal{V}),\omega<\infty$.
@@
[…]
@@
$\begin{aligned} h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})&=\lvert\eta_{\theta}^{\prime}(v_{\ddot{\tau}})\rvert\operatorname{\mathrm{N}}\left[\eta_{\theta}(v_{\ddot{\tau}});\eta_{\theta}(v_{\dot{\tau}}),(\ddot{\tau}-\dot{\tau})\rho_{\theta}^{2}(y_{\dot{\tau}})\right]\\ &\quad\times\exp\left\{\rho_{\theta}^{-2}(y_{\dot{\tau}})\{\Delta_{\theta}(\eta_{\theta}(v_{\ddot{\tau}}),y_{\dot{\tau}})-\Delta_{\theta}(\eta_{\theta}(v_{\dot{\tau}}),y_{\dot{\tau}})\}\right\}.\end{aligned}$ (9)
@@
[…]